Need ASAP!! Please help!!
2. Determine the coefficient of restitution for the bouncing of a ball from a hardwood foot 10 pts. as illustrated below. .9 m 7 m
Answer #1


for coefficient of restituition is CR = sqrt(h/H)

given that h = 0.7 m

and H = 0.8 m

then CR = sqrt(h/H) = sqrt(0.7/0.8) = 0.94